# BiSSL: Enhancing the Alignment Between Self-Supervised Pretraining and Downstream Fine-Tuning via Bilevel Optimization

### v4

@@ C.1 Computation Times
We report the training times required for both pretext pretraining and BiSSL under the experimental setup described in Section 4.2.1 and 4.2.2. SimCLR pretraining was conducted on 8x A100 GPUs, while BYOL used 4x A100 GPUs. BiSSL training was measured on the DTD dataset using 4x A40 GPUs. Since the number of gradient steps is kept constant across all datasets in the main experiments, this measurement is representative of BiSSL’s computational cost across other datasets as well.
The reported BiSSL training time reflects the total training time required on a single GPU, therefore we adjust for multi-GPU usage (e.g., for BiSSL, we multiply the total time taken by 4, as we used 4 GPUs). The results are documented in Table 6, and we observe that BiSSL requires only a small fraction of the computational cost of pretraining.
Table 6: Total GPU hours required for pretext pretraining and BiSSL training. BiSSL requires only fraction of the computational cost of pretext pretraining.
SimCLR BYOL
Pretext Training $1464$ GPU Hours $1380$ GPU Hours
BiSSL Training $68$ GPU Hours $70$ GPU Hours
C.2 Top-5 Classification Accuracies
Table 7 outlines the corresponding top-5 accuracies of the experiments described in Section 4.3, which re-emphasizes the performance improvements imposed by BiSSL as initially implied by Table 1. Likewise, Figure 4 outlines the corresponding top-5 accuracies of the experiments described in Section 4.3.1.
@@ C.2 Top-5 Classification Accuracies
Only FT $91.3\pm 0.1$ $100.0\pm 0.0$ $96.4\pm 0.1$ $75.2\pm 0.4$ $79.4\pm 0.4$ $93.7\pm 0.2$ $81.8\pm 0.5$ $85.8\pm 0.6$ $94.8\pm 0.4$ $98.1\pm 0.1$ $94.5\pm 0.2$
BiSSL+FT $\mathbf{93.1\pm 0.1}$ $100.0\pm 0.0$ $\mathbf{96.9\pm 0.1}$ $\mathbf{86.5\pm 0.2}$ $\mathbf{81.5\pm 0.1}$ $93.7\pm 0.1$ $\mathbf{85.0\pm 0.2}$ $\mathbf{88.7\pm 0.3}$ $\mathbf{96.5\pm 0.1}$ $\mathbf{98.7\pm 0.1}$ $\mathbf{95.8\pm 0.1}$
[…]
@@
Only FT $91.8\pm 0.1$ $100.0\pm 0.0$ $96.3\pm 0.1$ $77.5\pm 0.4$ $77.8\pm 0.3$ $93.9\pm 0.2$ $81.8\pm 0.6$ $84.9\pm 0.4$ $95.6\pm 0.2$ $98.0\pm 0.1$ $93.6\pm 0.3$
BiSSL+FT $91.8\pm 0.2$ $100.0\pm 0.0$ $\mathbf{96.7\pm 0.1}$ $\mathbf{86.7\pm 0.2}$ $\mathbf{80.1\pm 0.2}$ $94.1\pm 0.2$ $\mathbf{87.5\pm 0.3}$ $\mathbf{87.7\pm 0.3}$ $\mathbf{97.2\pm 0.1}$ $\mathbf{98.7\pm 0.1}$ $\mathbf{95.7\pm 0.1}$
[…]
Figure 4: Top-5 test classification accuracies on the Flowers dataset for separate models pretrained with SimCLR for different durations, comparing the conventional and BiSSL training pipelines. The corresponding top-1 accuracies are shown in Figure 3.
C.3 Impact of Upper-Level Iterations and the IJ Term
This section presents ablation studies showcasing how certain modifications in the BiSSL upper-level optimization affect downstream classification performance. We specifically modify the number of upper-level iterations and the inclusion of the implicit Jacobian (IJ) term in (2).
Due to resource constraints, we evaluate these variations using a representative subset of the 12 datasets from the main experiments (Section 4.1). Specifically, we chose the Pets [60], DTD [59], VOC07 [58], and Flowers [62] datasets, which together cover a diverse range of image domains. Pets and Flowers contain fine-grained object categories, VOC07 features multi-label natural scenes, and DTD consists of texture-based inputs.
All experiments in this section use SimCLR as the pretext task.
Upper-Level Iterations
@@ C.3 Impact of Upper-Level Iterations and the IJ Term
To evaluate this design choice, we conduct experiments with $N_{U}=1$ to match conventional bilevel update schedules, while keeping all other aspects identical to the default BiSSL setup in Section 4.2.2, aside from performing the subsequent fine-tuning hyperparameter grid search on 50 different hyperparameter combinations. Since $N_{U}=1$ results in fewer total upper-level updates, we also experiment with a longer training duration by increasing the number of training stage alternations from $T=500$ to $T=4000$, matching the total number of gradient updates conducted in the default setup. Notice however that this results in significantly more number of total lower-level steps ($20\cdot 4000=80,000$) conducted compared to the default setting ($40\cdot 500=20,000$). Despite this, the results in the two middle rows of Table 8 show that neither of the $N_{U}=1$ variants outperforms the default configuration in the top row with $N_{U}=8$ and $T=500$.
Discarding the IJ Term
We further examine the importance of the first term in the upper-level objective of Equation (2) that includes the IJ, which accounts for the influence of lower-level updates on the downstream loss. To do so, we discard this term and retain only the downstream loss $\mathcal{L}^{D}({\bf\it\theta}_{D},{\bf\it\phi}_{D})$ in the upper-level objective. The rest of the setup follows the default configuration in Section 4.2.2, except that the fine-tuning hyperparameter grid search was conducted for 50 different combinations. As shown in the final row of Table 8, removing the IJ term leads to notable drops in downstream accuracy, underscoring its role in guiding the BiSSL optimization toward better downstream task alignment.
@@
Table 8: Ablation results showing the impact on downstream classification accuracy from setting $N_{U}=1$ and from removing the first term of the upper-level objective in Equation (2). All experiments use SimCLR as the pretext task. The top row ($N_{U}=8,T=500$) corresponds to the default BiSSL configuration described in Section 4.2.2. The four datasets were selected based on input diversity and computational resource constraints.
BiSSL+FT Pets DTD VOC07 Flowers
$N_{U}=8,T=500$ $79.6\pm 0.3$ $63.4\pm 0.4$ $71.8\pm 0.1$ $85.6\pm 0.2$
$N_{U}=1,T=500$ $75.2\pm 0.4$ $61.5\pm 0.5$ $71.2\pm 0.1$ $82.2\pm 0.3$
$N_{U}=1,T=4000$ $77.6\pm 0.4$ $63.1\pm 0.3$ $71.2\pm 0.1$ $84.4\pm 0.2$
IJ Discarded $75.8\pm 0.4$ $62.9\pm 0.2$ $71.8\pm 0.1$ $82.0\pm 0.3$
C.4 Masked Autoencoder
Our main experiments demonstrate the effectiveness of BiSSL using contrastive pretext tasks and ResNet-based architectures. We here evaluate BiSSL in a slightly different setting using a Masked Autoencoder (MAE) [7].
Table 9: Downstream classification performance when using a MAE as the pretext task. BiSSL improves performance and reduces variance across all four datasets. The four datasets were selected based on input diversity and computational resource constraints.
MAE Pets DTD VOC07 Flowers
Only FT $81.3\pm 0.5$ $61.3\pm 0.7$ $69.5\pm 0.3$ $81.1\pm 1.1$
BiSSL+FT $\mathbf{83.2\pm 0.2}$ $\mathbf{63.5\pm 0.4}$ $\mathbf{71.8\pm 0.2}$ $\mathbf{87.2\pm 0.3}$
[…]
We evaluate performance on a representative subset of the twelve downstream datasets used in the main experiments, due to resource limitations: Pets [60], DTD [59], VOC07 [58], and Flowers [62]. These datasets were selected to represent a broad range of data distributions and classification challenges, from fine-grained object categories (Pets, Flowers), to texture-based recognition (DTD), and multi-label natural scenes (VOC07).
The results in Table 9 show that BiSSL again yields consistent improvements across all datasets and reduces variance in downstream accuracy compared to solely conducting conventional fine-tuning.
C.5 Additional Baseline Comparisons
This section introduces two additional baseline methods that involve intermediate modifications to the pretrained backbone prior to fine-tuning, aiming to provide further context for the effectiveness of BiSSL.
Due to computational constraints, we evaluate these baselines on a subset of four datasets from the main experiments, namely the Pets [60], DTD [59], VOC07 [58], and Flowers [62] datasets. These datasets were selected to represent distinct challenges and input domains in visual recognition, including fine-grained classification (Pets, Flowers), texture-based classification (DTD), and multi-label object recognition in natural scenes (VOC07).
In all experiments within this section, we use SimCLR as the pretext task.
Accuracy $w=\mathbf{0}$ $w=\mathbf{0.05}$ $w=\mathbf{0.1}$ $w=\mathbf{0.25}$ $w=\mathbf{0.5}$ $w=\mathbf{0.75}$ Only FT BiSSL+FT
Top-1 $60.1\pm 0.3$ $61.6\pm 0.2$ $61.7\pm 0.2$ $\underline{62.2\pm 0.3}$ $60.8\pm 0.3$ $60.3\pm 0.5$ $60.3\pm 0.9$ $\mathbf{63.4\pm 0.3}$
Top-5 $\underline{86.3\pm 0.3}$ $\underline{86.2\pm 0.3}$ $\underline{86.2\pm 0.3}$ $\underline{86.1\pm 0.3}$ $\underline{86.2\pm 0.3}$ $85.7\pm 0.3$ $85.8\pm 0.6$ $\mathbf{88.7\pm 0.3}$
Table 11: Comparison of downstream classification accuracy using alternative baseline methods: a weighted combination of pretext and downstream objectives, and NoisyTune. All experiments use SimCLR as the pretext task. The four datasets were selected based on input diversity and computational resource constraints. BiSSL significantly outperforms all baselines.
Pets DTD VOC07 Flowers
Only FT $73.2\pm 0.3$ $60.3\pm 0.9$ $71.0\pm 0.1$ $82.6\pm 0.3$
PretextDownstreamSum+FT $76.4\pm 0.4$ $62.2\pm 0.3$ $67.1\pm 0.1$ $83.8\pm 0.1$
NoisyTune+FT $73.4\pm 0.3$ $60.0\pm 0.7$ $70.3\pm 0.2$ $82.6\pm 0.2$
BiSSL+FT  $\mathbf{79.6\pm 0.3}$ $\mathbf{63.4\pm 0.4}$ $\mathbf{71.8\pm 0.1}$ $\mathbf{85.6\pm 0.2}$
Weighted Sum of Pretext and Downstream Objectives
The first baseline attempts a simple alignment of the pretext and downstream tasks by solving the single-level optimization problem:
We use the lower-level optimizer configuration from BiSSL (see Section 4.2.2) and train for $4000$ steps, matching the total number of upper-level iterations in the default BiSSL configuration. Fine-tuning follows the main setup but with a grid search over 50 hyperparameter combinations. Backbone gradients for the two loss terms are computed separately on their respective distinct mini-batches, scaled by $1-w$ and $w$ respectively, and then summed prior to the update step.
To determine the suitable size of $w$, we conducted a sweep on the DTD dataset over different values of $w$ in the range between 00 and $1$, evaluating both top-1 and top-5 classification performance. For each value of $w$, we conduct the subsequent fine-tuning hyperparameter grid search over 25 different combinations. The results of the sweep are shown in Table 10.
The results imply that $w=0.25$ provides the best top-1 classification accuracy. In contrast to BiSSL, the top-5 accuracy does generally not show a significant increase compared to conventional fine-tuning, regardless of the value of $w$. Setting $w=0$ essentially reduces the training to pretext-only learning, reflected by comparable performance to conventional fine-tuning (the "Only FT" column), with slightly lower top-1 accuracy and slightly higher top-5 accuracy. Based on the results, we consider $w=0.25$ as the optimal value for this baseline.
Using $w=0.25$, we summarize the baseline experiments in the "PretextDownstreamSum" row in Table 11, where BiSSL significantly outperforms this baseline across all four datasets.
NoisyTune
We also evaluate NoisyTune [44], which applies Gaussian noise to the backbone weights prior to fine-tuning. We use a noise scale of $0.015$, and conduct a hyperparameter grid search over $50$ different combinations. Results are included in the "NoisyTune" row of Table 11, where BiSSL again demonstrates superior accuracy across all datasets.
C.6 Varying the Downstream Dataset Size
Our downstream datasets already vary widely in size, ranging from $1,020$ to $60,600$ samples, with Table 1 showing no clear correlation between dataset size and accuracy. To further investigate whether this trend persists when considering a single dataset, we conducted additional experiments on the Pets [60] dataset, varying the number of samples per class. The SimCLR pretext task [1] was used, and the fine-tuning hyperparameter grid search was conducted over $25$ combinations, while all other settings were maintained as in the original setup described in Section 4.2.2. We evaluated accuracy using $5$, $10$, $25$, $50$, $75$, samples per class.
Table 12: Varying the Number of Samples per Class on the Pets Dataset. SimCLR is used as the pretext task.
Samples Per Class $\mathbf{5}$ $\mathbf{10}$ $\mathbf{25}$ $\mathbf{50}$ $\mathbf{75}$
Only FT $33.4\pm 0.5$ $45.4\pm 0.5$ $62.5\pm 0.6$ $70.5\pm 0.5$ $73.1\pm 0.2$
BiSSL+FT $\mathbf{37.2\pm 0.6}$ $\mathbf{52.3\pm 0.3}$ $\mathbf{67.2\pm 0.2}$ $\mathbf{73.0\pm 0.3}$ $\mathbf{77.7\pm 0.1}$
[…]
Results are presented in Table 12, which show that BiSSL consistently outperforms conventional fine-tuning, regardless of the number of samples available per class.
C.7 Fine-Tuning on ImageNet
@@
We expect the gains from BiSSL on ImageNet classification [36] as a downstream task to be limited in this specific setting, as pretraining is also conducted on ImageNet, which minimizes the distribution discrepancy between the pretext and downstream tasks. Nonetheless, for the sake of transparency, we conducted fine-tuning experiments on ImageNet within the bounds of our available computational resources.
@@
Our setup consists of $50$ fine-tuning epochs, using a hyperparameter grid search over $25$ configurations. For BiSSL, we use $T=100$ training stage alternations, while keeping the rest of the experimental setup consistent with the description in Section 4.2.2. We use SimCLR as the pretext task. Following standard practice in SSL literature for ImageNet fine-tuning [1, 4, 2], we report the top-1 validation accuracy on models trained with the $1\%$ and $10\%$ subsets using the official splits from [1]. Additionally, we also include results on the full ImageNet dataset ($100\%$). Table 13 shows the results of these experiments.
@@
Table 13: Top-1 validation classification accuracy on ImageNet using $1\%$, $10\%$, and $100\%$ of the labeled training data. SimCLR is used as the pretext task.
@@
Only FT $29.7\pm 0.1$ $50.9\pm 0.1$ $68.2\pm 0.1$
BiSSL+FT $\mathbf{31.8\pm 0.1}$ $51.1\pm 0.1$ $68.2\pm 0.1$
[…]
@@
As expected, BiSSL yields only modest improvements on ImageNet. The most notable gain appears in the $1\%$ setting, which may be attributed to a greater distribution mismatch between the small labeled subset and the pretraining dataset. This represents the type of scenario where BiSSL is particularly effective at improving alignment. In contrast, the $10\%$ subset is more representative of the full dataset, and the relative advantage of BiSSL correspondingly diminishes.
C.8 Visual inspection of Latent Features
During the evaluation, it is important to note that the batch normalization layers [63] of the pretrained backbones utilize the running mean and variance inferred during training. Since these pretrained backbones have not been exposed to the downstream datasets during training, their batch normalization statistics may not be optimal for these new datasets. To address this, the downstream training dataset is divided into batches of $256$ samples, and roughly $100$ batches are then forward-passed through the backbones. This procedure ensures that the batch normalization statistics are better suited to the downstream datasets, thereby providing a fairer comparison of the learned representations.
For the dimensionality reduction and visualization of these latent features, the t-Distributed Stochastic Neighbor Embedding (t-SNE) [69] technique is employed. This method allows us to visually assess the clustering and separation of features in the latent space, providing qualitative insights into the semantic structure of the representations learned through BiSSL.
Figures 3 and 5 to 10 illustrate the outcomes of these visual inspections on a selection of the downstream datasets described in Section 4.1, highlighting the differences in feature representations between conventional pretext pretraining and BiSSL.
Figure 5: Features from lower-level backbones after applying BiSSL (left) or pretext pretraining (right) on the CIFAR10 [53] dataset.
Figure 6: Features from lower-level backbones after applying BiSSL (left) or pretext pretraining (right) on the CUB200 [54] dataset.
Figure 7: Features from lower-level backbones after applying BiSSL (left) or pretext pretraining (right) on the Caltech-101 [61] dataset.

### v5

@@ C.1 Computation Times
We report the training times required for both pretext pretraining and BiSSL under the experimental setup described in Section 4.2.1 and 4.2.2. SimCLR pretraining was conducted on 8x A100 GPUs, and BiSSL training was measured on the DTD dataset using 4x A40 GPUs. Since the number of gradient steps is kept constant across all datasets in the main experiments, this measurement is representative of BiSSL’s computational cost across other datasets as well. For transparency, we also report fine-tuning training times. The duration naturally varies with dataset size, and we present the range of computation times across the datasets listed in Table 1, measured on a single A40 GPU.
The reported training times reflects the total training time required on a single GPU, therefore we adjust for multi-GPU usage (e.g., for BiSSL, we multiply the total time taken by 4, as we used 4 GPUs). The results are documented in Table 10. BiSSL incurs some additional overhead over fine-tuning, but requires only a small fraction of the computational cost of pretraining.
Table 10: Total GPU hours required for pretext pretraining and BiSSL training with the SimCLR pretext task. BiSSL introduces some overhead relative to fine-tuning, while its computational cost is only a fraction of that required for pretext pretraining.
Computation Time
Pretext Training $1464$ GPU Hours
BiSSL Training $83$ GPU Hours
Fine-Tuning 11-99 GPU Hours
C.1.1 Remarks on Hessian Vector Products
A significant portion of the computational overhead in BiSSL arises from calculating Hessian-vector products (HVPs) during each upper-level gradient update. For each upper-level gradient update step, the upper-level compute $N_{c}$ HVPs (five in the experiments of this paper) using the CG algorithm. Each HVP requires an additional backward pass, effectively adding $2N_{c}$ backward passes. Memory usage is also increased, as activations must be stored for the second backward pass.
The memory and runtime costs depend on the HVP implementation, typically requiring two to three times the memory of standard gradient computations and taking between two and four times longer to compute (Dagréou et al., May 7, 2024). Further multiplying the runtime values by $N_{c}$ highlights the substantial additional compute required by upper-level optimization relative to standard gradient computation. In principle, memory and runtime are expected to scale roughly linearly with backbone size. Under heavy memory load, however, additional runtime overhead can occur, potentially resulting in super-linear scaling for very large models. The latter could however be mitigated by combining BiSSL with parameter-efficient fine-tuning (PEFT) techniques, such as LoRA Hu et al. (2022), on both the upper and lower-levels, which we leave for future work.
@@ C.1.1 Remarks on Hessian Vector Products
FT $91.3\pm 0.1$ $100.0\pm 0.0$ $96.4\pm 0.1$ $75.2\pm 0.4$ $79.5\pm 0.4$ $93.7\pm 0.2$ $81.9\pm 0.5$ $85.8\pm 0.6$ $94.2\pm 0.4$ $98.1\pm 0.1$ $94.5\pm 0.2$
BiSSL+FT $\mathbf{93.4\pm 0.0}$ $100.0\pm 0.0$ $\mathbf{96.6\pm 0.2}$ $\mathbf{84.2\pm 0.2}$ $\mathbf{81.2\pm 0.2}$ $93.7\pm 0.2$ $\mathbf{84.3\pm 0.3}$ $\mathbf{87.8\pm 0.3}$ $\mathbf{96.1\pm 0.1}$ $\mathbf{98.6\pm 0.1}$ $\mathbf{95.3\pm 0.2}$
Avg Diff $\mathbf{+2.1}$ $0.0$ $\mathbf{+0.2}$ $\mathbf{+9.0}$ $\mathbf{+1.7}$ $0.0$ $\mathbf{+2.4}$ $\mathbf{+2.0}$ $\mathbf{+1.9}$ $\mathbf{+0.5}$ $\mathbf{+0.8}$
@@
FT $91.9\pm 0.1$ $100.0\pm 0.0$ $96.3\pm 0.1$ $7.5\pm 0.4 77$$7.8\pm 0.3$ $93.9\pm 0.2$ $81.8\pm 0.6$ $84.9\pm 0.5$ $95.6\pm 0.2$ $98.0\pm 0.1$ $93.6\pm 0.3$
BiSSL+FT $\mathbf{92.3\pm 0.1}$ $100.0\pm 0.0$ $96.4\pm 0.1$ $\mathbf{84.5\pm 0.2}$ $\mathbf{79.7\pm 0.1}$ $94.0\pm 0.2$ $\mathbf{86.2\pm 0.3}$ $\mathbf{86.7\pm 0.4}$ $\mathbf{96.6\pm 0.2}$ $\mathbf{98.5\pm 0.1}$ $\mathbf{94.2\pm 0.1}$
Avg Diff $\mathbf{+1.3}$ $0.0$ +0.1+0.1 $\mathbf{+7.0}$ $\mathbf{+1.9}$ +0.1+0.1 $\mathbf{+4.4}$ $\mathbf{+1.8}$ $\mathbf{+1.0}$ $\mathbf{+0.5}$ $\mathbf{+0.6}$
C.2 Top-5 Classification Accuracies
Table 11 outlines the corresponding top-5 accuracies of the experiments described in Section 4.3, which re-emphasizes the performance improvements imposed by BiSSL as initially implied by Table 1.
C.3 Impact of Varying Upper-Level Iterations
@@
Experiments were conducted using the same five downstream datasets used in Section 4.4.1 namely Pets, DTD, VOC07, Flowers and CUB200. As discussed there, these datasets span both coarse- and fine-grained classification tasks across diverse and largely disjoint domains. Also consistent with that setup, we use SimCLR as the pretext task.
@@
We conduct experiments with $N_{U}=1$ to match conventional bilevel update schedules, while keeping all other aspects identical to the default BiSSL setup in Section 4.2.2, aside from performing the subsequent fine-tuning hyperparameter grid search on 50 different hyperparameter combinations. Since $N_{U}=1$ results in fewer total upper-level updates, we also experiment with a longer training duration by increasing the number of training stage alternations from $T=500$ to $T=4000$, matching the total number of gradient updates conducted in the default setup. Notice however that this results in significantly more number of total lower-level steps ($20\cdot 4000=80,000$) conducted compared to the default setting ($40\cdot 500=20,000$). Despite this, the results in Table 12 show that neither of the $N_{U}=1$ variants provide any clear general advantage over the default configuration in the bottom row with $N_{U}=8$ and $T=500$.
@@
Table 12: Impact on downstream classification accuracy from setting $N_{U}=1$. All experiments use SimCLR as the pretext task. The bottom row ($N_{U}=8,T=500$) corresponds to the default BiSSL configuration described in Section 4.2.2.
BiSSL+FT Pets DTD VOC07 Flowers CUB200
$N_{U}=1,T=500$ $76.1\pm 0.5$ $61.8\pm 0.4$ $71.0\pm 0.1$ $82.4\pm 0.5$ $50.0\pm 0.9$
$N_{U}=1,T=4000$ $78.0\pm 0.2$ $63.2\pm 0.3$ $71.4\pm 0.1$ $84.4\pm 0.2$ $56.0\pm 0.4$
$N_{U}=8,T=500$ $78.3\pm 0.2$ $64.2\pm 0.4$ $71.5\pm 0.1$ $84.1\pm 0.2$ $55.7\pm 0.2$
C.4 Impact of Varying the Lower-Level Regularization Weight
As discussed in Section 3.3, the magnitude of the lower-level regularization weight $\lambda$ controls the extent to which the upper-level downstream objective influences the lower-level pretext optimization. All main experiments in this paper use a constant value of $\lambda=0.001$. To examine the effect of this parameter more systematically, we repeat the default experimental setup described in Section 4.2 while varying $\lambda$. For each value, we run the full fine-tuning hyperparameter search spanning $24$ configurations. We use SimCLR as the pretext task and report results on the DTD dataset in Figure 5.
Figure 5: Top-1 test accuracies on the DTD dataset for models pretrained with SimCLR across different values of $\lambda$ used in the BiSSL training stage. When $\lambda$ is sufficiently large, BiSSL consistently yields substantial performance gains, with additional improvements possible through tuning around $\lambda=0.001$.
The results show that BiSSL provides notable improvements whenever $\lambda$ is large enough for the upper-level signal to meaningfully shape the lower-level optimization. More precise tuning yields further gains, with $\lambda=0.001$ achieving the highest accuracy among the tested values. This outcome aligns with the interpretation in Section 3.3. When $\lambda$ is too small, the lower-level effectively behaves like continued pretext pretraining, providing little alignment with the downstream objective. At the same time, the theoretical setup in Section A.1 requires $\lambda$ to be sufficiently large in order to preserve lower-level convexity. Values that are too small can violate these assumptions, which may also help explain the negligible or absent performance gains observed in that regime.
C.5 Fine-Tuning on ImageNet
@@
We expect the gains from BiSSL on ImageNet classification (Deng et al., 2009) as a downstream task to be limited in this specific setting, as pretraining is also conducted on ImageNet, which minimizes the distribution discrepancy between the pretext and downstream tasks. Nonetheless, for the sake of transparency, we conducted fine-tuning experiments on ImageNet within the bounds of our available computational resources.
@@
Our setup consists of $50$ fine-tuning epochs, using a hyperparameter grid search over $25$ configurations. For BiSSL, we use $T=100$ training stage alternations, while keeping the rest of the experimental setup consistent with the description in Section 4.2.2. We use SimCLR as the pretext task. Following standard practice in SSL literature for ImageNet fine-tuning (Chen et al., 2020; Grill et al., 2020; Bardes et al., 2022), we report the top-1 validation accuracy on models trained with the $1\%$ and $10\%$ subsets using the official splits from Chen et al. (2020). Additionally, we also include results on the full ImageNet dataset ($100\%$).
@@
Table 13: Top-1 validation classification accuracy on ImageNet using $1\%$, $10\%$, and $100\%$ of the labeled training data. SimCLR is used as the pretext task. Accuracies statistically significantly higher from their counterparts are marked in the bold.
@@
FT $29.7\pm 0.1$ $50.9\pm 0.1$ $68.2\pm 0.1$
BiSSL+FT $\mathbf{31.5\pm 0.1}$ $\mathbf{51.1\pm 0.1}$ $68.3\pm 0.1$
Avg Diff $\mathbf{+1.8}$ $\mathbf{+0.2}$ +0.1+0.1
@@
Table 13 shows the results of these experiments. The most notable gain in the $1\%$ setting may be attributed to a greater distribution mismatch between the small labeled subset and the pretraining dataset, representing the type of scenario where BiSSL is particularly effective at improving alignment. In contrast, the $10\%$ subset is more representative of the full dataset, and the relative advantage of BiSSL correspondingly decrease.
C.6 Larger Image Size with Official SimCLR Backbone Weights
To assess how BiSSL scales to higher image resolutions and to verify its compatibility with off-the-shelf pretrained models, we benchmark BiSSL using the official SimCLR ResNet-50 backbone weights released by Chen et al. (2020), trained with an input size of $224\crossproduct 244$.
Table 14: Downstream classification performance using 224px images with the official pretrained backbone via SimCLR. BiSSL improves accuracy with statistical significance across four out of the five datasets.
SimCLR 224px Pets DTD VOC07 Flowers CUB200
FT $87.0\pm 0.3$ $69.5\pm 0.4$ $75.1\pm 0.1$ $89.7\pm 0.3$ $66.0\pm 0.2$
BiSSL+FT $\mathbf{87.7\pm 0.2}$ $\mathbf{70.7\pm 0.3}$ $\mathbf{77.2\pm 0.1}$ $89.8\pm 0.3$ $\mathbf{71.6\pm 0.3}$
Avg Diff $\mathbf{+0.7}$ $\mathbf{+1.2}$ $\mathbf{+2.1}$ +0.1+0.1 $\mathbf{+5.6}$
Because the official SimCLR checkpoints do not include their original projection head, we trained a replacement projection head with a similar architecture on the frozen backbone for $20$ epochs using the SimCLR objective. To accommodate computational limits, we used a batch size of $1024$ instead of the original $4096$. For BiSSL, we further reduced total training time by setting $T=200$, $N_{U}=4$, and $N_{L}=10$ (see Algorithm 1 for details). The subsequent fine-tuning hyperparameter grid search is conducted over $25$ different combinations. We used the larger input resolution of $244\crossproduct 244$, with all other settings identical to those described in Sections 4.2.1 and 4.2.2. Experiments were performed on the same five downstream classification dataset used in Section 4.4.1, namely Pets, DTD, VOC07, Flowers and CUB200. As argued in that section, these datasets cover both coarse- and fine-grained classification tasks across diverse and mostly disjoint domains.
While the setup does not replicate the full-scale SimCLR configuration (lacking the original projection, batch size, and employing a shorter BiSSL training duration), it nonetheless provides a meaningful indicator of BiSSL’s robustness. As shown in Table 14, BiSSL sustains its ability to significantly enhance downstream performance, even under these conditions.
C.7 Visual inspection of Latent Features
Figure 4 illustrated features on the flowers dataset from SimCLR-pretrained backbones to BiSSL-trained backbones. Figures 6 to 11 illustrate similar backbone features on another selection of the downstream datasets described in Section 4.1. The trend, where BiSSL appear to better align the backbone with the downstream task, persist even for datasets where the accuracy gains from applying BiSSL are minimal or absent.